OpenStudy (anonymous):

solve 25^(x+2)=125^(x)

OpenStudy (turingtest):

write 125 as a power of 25

OpenStudy (anonymous):

\[\large 25=5^2 \\ \large 125=5^3\]

OpenStudy (anonymous):

\[ 25^{x+2}=125^x\\ 25^{x+2}=(25^{3/2})^x\\ 25^{x+2}=25^{3x/2}\\ x+2=3x/2 \]Can you solve from there?

OpenStudy (turingtest):

lol I mean power of 5

OpenStudy (turingtest):

I guess both work :P

OpenStudy (anonymous):

Yep, both work. Writing as a power of five will be easier for you if you are not yet familiar with fractions as exponents.

OpenStudy (anonymous):

nbouscal. is it solving like a regular equation after that?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

one minute.

OpenStudy (anonymous):

is it x=1? 2=1.5-x 2=.5x 1=x

OpenStudy (anonymous):

No.\[ x+2=\frac{3x}2\\ 2(x+2)=3x \]

OpenStudy (anonymous):

so x=4 (after dispersing the 2) 2x+4=3x 4=x

OpenStudy (anonymous):

Right. You can plug that back in and check to see that it's right. Use your calculator, you'll see that \(25^6=125^4.\)

OpenStudy (anonymous):

yay!!! Thank you!!!!